OpenStudy (anonymous):

Four more than twice a number is seven more than the number. What is the number

OpenStudy (ja1):

@gabriel4031 do you need the equation or solve it?

OpenStudy (anonymous):

Let the number be x therefore 2x+4=x+7 2x-x=7-4 x=3 so the required number is 3

OpenStudy (anonymous):

A. –4 B. 9.6 C. 8 D. 16

OpenStudy (anonymous):

solve

OpenStudy (anonymous):

question?????